# CICLe: Conformal In-Context Learning for Largescale Multi-Class Food Risk Classification

### v2

@@ 5.2 Experimental Results
Table 1: Macro-averaged $F_{1}$-score (mean & maximum) over 5 CV splits. Blue cells highlight the best mean score per column and classification task. For hazard and product scores are only evaluated on well supported classes.
@@ 5.2.1 Traditional Baselines
The $F_{1}$-scores on classification (overall, as well as $\mathcal{C}_{high}$ and $\mathcal{C}_{low}$ introduced in Figure 2) are presented in Table 1. All the classifiers outperform both naive baselines, except for the $\mathcal{C}_{high}$ of hazard-category. In this task, $\mathcal{C}_{high}$ only consists of the majority class. Since the BOW-based classifiers in general perform worse than TF-IDF, we only report them in the Appendix (see Table 8).
@@ 5.2.2 Transformers
Interestingly, RoBERTa and XLM-R both achieve lower and less stable $F_{1}$-score for product and product-category. This is not seen in other classifiers. By inspecting the data, we see that in both of these tasks brand names (e.g. "SnoBalls", "Sriracha") or hard-to-define word creations (e.g. "pecan caramel stars", "Vanilla Cups") make up the relevant part for classification. These are hard to learn during fine-tuning, as they are both very specific and most probably underrepresented in the data used for pre-training. Analysis of the spans in product-title shows that TF-IDF-LR often infers the product from the supplier and/or the hazard in such cases.
@@ 5.2.3 Large Language Models
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: CP performance on the best traditional classifier and the best encoder-only Transformer. The optimal prediction set (red dashed line) would have one element which is the true class with a probability of 1
@@
Focusing on GPT-CICLe, Table 2 shows that it outperforms its base classifier TF-IDF-LR on all tasks as well as all reduced prompting baselines for three out of four tasks. For higher numbers of classes, the performance of GPT-CICLe drops in $\mathcal{C}_{low}$ compared to the other prompts (see Table 9). We attribute this to the higher $\alpha$ in these cases: here we sacrifice CP-guarantees for faster prediction. GPT-CICLe also produces the lowest prompt lengths at the same $F_{1}$ (Table 2). Here, we argue, that the base classifier used by GPT-MAX-$k$ and GPT-CICLe is able to produce more concise contexts for the LLM than simple cosine similarity. While this is already a step towards energy-efficient prompting, CICLe can avoid accessing the LLM completely in some cases: the extent depends on the task and the $\alpha$, but ranges between accessing the LLM in only $60\%$ to $98\%$ of cases in our tasks.
@@ 7 Conclusions
We present the first dataset for large-scale multi-class classification of short texts describing food recalls, which we release publicly under a CC BY-NC-SA 4.0 license. The dataset contains labels annotated by experts on food products and hazards on two levels of granularity, coarse (tens) and fine (hundreds). We carry out benchmarking, analysing the performance of naive, traditional, and deep-learning classifiers for all four classification tasks in the proposed dataset, showing the potential of traditional classifiers on this dataset. By employing this classifier, we also experimented with prompting, proposing a novel LLM-in-the-loop framework that leverages Conformal Prediction to establish more accurate and energy-efficient performance.
@@ Appendix A Data sample and statistics:
In the following Tables and Figures, we provide additional insights into the dataset shared with this publication. Table 3 lists the original websites our texts were collected from, Table 4 provides a statistical overview over the class- and language-distribution in the data, and Figure 5 is a heatmap highlighting co-occurrence of classes in the hazard-category and product-category tasks. Finally, we share a few data samples along with their labels in Figure 6.

### v3

@@ 5.2 Experimental Results
Table 1: Macro-averaged $F_{1}$-score (mean & maximum) over 5 CV splits. Blue cells highlight the best mean score per column and classification task. For hazard and product scores are only evaluated on well-supported classes. The grayed scores are added for completeness, but not comparable, as $\mathcal{C}_{high}$ of hazard-category consists of only one class.
@@ 5.2.1 Traditional Baselines
The $F_{1}$-scores on classification (overall, as well as $\mathcal{C}_{high}$ and $\mathcal{C}_{low}$ introduced in Figure 2) are presented in Table 1. All the classifiers outperform both naive baselines, except for the $\mathcal{C}_{high}$ of hazard-category. In this task, $\mathcal{C}_{high}$ only consists of the majority class. This means that classification on this task is not useful as there is only one label present in the test set. Every class not predicted to this label (even if correctly predicted in general) will therefore diminish the macro $F_{1}$. Since the BOW-based classifiers in general perform worse than TF-IDF, we only report them in the Appendix (see Table 8).
@@ 5.2.2 Transformers
Interestingly, RoBERTa and XLM-R both achieve lower and less stable $F_{1}$-score for product and product-category. This is not seen in other classifiers. By inspecting the data, we see that in both of these tasks brand names (e.g. “SnoBalls”, “Sriracha”) or hard-to-define word creations (e.g. “pecan caramel stars”, “Vanilla Cups”) make up the relevant part for classification. These are hard to learn during fine-tuning, as they are both very specific and most probably underrepresented in the data used for pre-training. Analysis of the spans in product-title shows that TF-IDF-LR often infers the product from the supplier and/or the hazard in such cases.
@@ 5.2.3 Large Language Models
Figure 4: CP performance on the best traditional classifier (green dashed line) and the best encoder-only Transformer (blue dashed line). The optimal prediction set (red dashed line) would have one element, which is the true class with a probability of 1.
@@
Focusing on GPT-CICLe, Table 2 shows that it outperforms its base classifier TF-IDF-LR on all tasks as well as all reduced prompting baselines for three out of four tasks. For higher numbers of classes, the performance of GPT-CICLe drops in $\mathcal{C}_{low}$ compared to the other prompts (see Table 9). We attribute this to the higher $\alpha$ in these cases: here we sacrifice CP guarantees for faster prediction. GPT-CICLe also produces the lowest prompt lengths at the same $F_{1}$ (Table 2). Here, we argue, that the base classifier used by GPT-MAX-$k$ and GPT-CICLe is able to produce more concise contexts for the LLM than simple cosine similarity. While this is already a step towards energy-efficient prompting, CICLe can avoid accessing the LLM completely in some cases: the extent depends on the task and the $\alpha$, but ranges between accessing the LLM in only $60\%$ to $98\%$ of cases in our tasks. The selection of alpha can in theory be done through an extensive hyperparameter search. As this becomes very computationally expensive for LLMs, we select $\alpha$ based on Figure 4 in our experiments (see Appendix C). We leave computationally efficient optimization of $\alpha$ for future research.
@@ 7 Conclusions
We present the first dataset for large-scale multi-class classification of short texts describing food recalls, which we release publicly under a CC BY-NC-SA 4.0 license. The dataset contains labels annotated by experts on food products and hazards on two levels of granularity, coarse (tens) and fine (hundreds).
Answering our research question R1, we find that the titles of food recall announcements mostly hold sufficient information for classification, although not all of them include information on both product and hazards (R1.1). A strong indication for this statement is that we are able to reach above random performance with multiple classifiers: we carry out benchmarking, analysing the performance of naive, traditional, and deep-learning classifiers for all four classification tasks in the proposed dataset, showing the potential of traditional classifiers on this dataset. Our results further indicate that products are harder to classify than hazards, which we attribute to the often creative choice of product names (R1.2).
Moving on to R2, we also experimented with prompting using one few-shot example per class. Here we see that gpt-3.5 performs comparably or better than the best non-LLM classifiers (R2.1). We also show that resource-efficient prompting at comparable predictive performance is possible by proposing a novel LLM-in-the-loop framework that leverages Conformal Prediction (R2.2).
@@ Acknowledgements
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
@@ Appendix A Data sample and statistics:
In this appendix, we provide additional insights into the dataset shared with this publication. Table 3 lists the original websites our texts were collected from, Table 4 provides a statistical overview over the class- and language-distribution in the data, and Figure 5 is a heatmap highlighting co-occurrence of classes in the hazard-category and product-category tasks. Finally, we share a few data samples along with their labels in Figure 6.